tahyla:

Sylvia is 12 years older than her sister Stacy. She is also 5 times as old as Stacy. how old is sylvia and stacy??

Sackrunnaz:

60

tahyla:

huh

Sackrunnaz:

oh nvm didnt read it all the way

tahyla:

kk

ummmmmmmigyourok:

1000

toga:

start by setting up some equations based on the information given: Let's say that Sylvia's current age is "S" and Stacy's current age is "T". From the first piece of information, we know that: S = T + 12 And from the second piece of information, we know that: S = 5T Now we can use substitution to solve for T: 5T = T + 12 4T = 12 T = 3 So Stacy is currently 3 years old. We can use this information to solve for Sylvia's age: S = 5T S = 5(3) S = so Sylvia is how old